# Unlocking the Non-Native Language Context Limitation: Native Language Prompting Facilitates Knowledge Elicitation

### v1

@@ Abstract
Multilingual large language models (MLLMs) struggle to answer questions posed in non-dominant languages, even though they have already acquired the relevant knowledge from their dominant language corpus. In contrast, human multilinguals can overcome this issue by invoking the relatively rich knowledge acquired from native language texts through Positive Native Language Transfer (PNLT). Inspired by this, we analogize the dominant language of MLLMs to the native language of human multilinguals, and propose Native Language Prompting (NatLan) to simulate the PNLT observed in human multilinguals. It explicitly creates native language contexts for MLLMs to facilitate the elicitation of the rich native language knowledge during question-answering, unlocking the limitations imposed by non-native language contexts on the effective application of knowledge. By employing multi-MLLM collaboration, NatLan reduces the workload on each MLLM in simulating PNLT and refines semantic transfer. On the C-Eval benchmark, NatLan provides up to a 10.1% average accuracy improvement and up to a 5.0% increase in the hard-level subset across five MLLMs, surpassing all top-notch related methods. Our code is available at https://github.com/AnonyNLP/NatLan.
@@ 1 Introduction
Figure 1: To address the failure of knowledge elicitation when directly answering in non-native language (Left), NatLan simulates the Positive Native Language Transfer of a human multilingual by utilizing two different MLLMs (Right). English meanings appear in <>absent<\ >< >.
Multilingual large language models (MLLMs) Brown et al. (2020); Achiam et al. (2023) have propelled the advancement of nearly all natural language processing tasks across various languages Xu et al. (2024); Li et al. (2024); Thakur et al. (2024). However, it’s observed that MLLMs perform poorly on questions articulated in non-dominant languages, as depicted in Figure 1 (Left), failing to answer some questions that they could address when presented in their dominant language (i.e., the language with the highest proportion during training, such as English for Llama Touvron et al. (2023a), which accounts for over 70% of the tokens in the pretraining corpus, significantly more than any other languages). This indicates that MLLMs are unable to effectively apply the rich knowledge acquired from dominant language contexts when operating in non-dominant language contexts. Existing work attributes this to the differing volumes Xue et al. (2021); ImaniGooghari et al. (2023) and quality Sitaram et al. (2023) of training data in various languages, which leads to insufficient training in non-dominant languages.
@@ 5 Experiments
In the proposed NatLan, the Transferor must be capable of translating the content from the target language (Chinese) into the native language (English) as accurately and coherently as possible. Therefore, we selected the Qwen series MLLMs Bai et al. (2023) as Transferors, for their leading capabilities in Chinese comprehension among all MLLMs. We chose Qwen models with 4B, 7B, and 14B parameters to analyze the effects of Transferors with varying capabilities on NatLan in §5.5 and §5.6. Additionally, we selected a five representative MLLMs with strong English comprehension skills and the capacity to understand Chinese to serve as Speakers. These include models from the Phi Abdin et al. (2024), Gemma Team et al. (2024), Mistral Jiang et al. (2023), and Llama Touvron et al. (2023b) series. For ease of joint deployment with the Transferor LLMs, all these Speaker LLMs possess a moderate parameter scale, ranging from 3.8B to 7B.
@@ A.2 Comparative Analysis of Chinese-to-English Translation Cases
The impact of the capital relocation on Palangkaraya is________\_\_\_\__ _ _ _. A. It is conducive to alleviating housing shortages B. It is conducive to alleviating traffic congestion C. It is conducive to the control of environmental pollution D. The improvement of infrastructure

### v2

@@ Abstract
Multilingual large language models (MLLMs) struggle to answer questions posed in non-dominant languages, even though they have acquired the relevant knowledge from their dominant language corpus. In contrast, human multilinguals can overcome such non-native language context limitations through Positive Native Language Transfer (PNLT). Inspired by the process of PNLT, we analogize the dominant language of MLLMs to the native language of human multilinguals, and propose Native Language Prompting (NatLan) to simulate the PNLT observed in human multilinguals. It explicitly creates native language contexts for MLLMs to facilitate the elicitation of the rich native language knowledge during question-answering, unlocking the limitations imposed by non-native language contexts. By employing multi-MLLM collaboration, NatLan reduces the workload on each MLLM in simulating PNLT and refines semantic transfer. On the C-Eval benchmark, NatLan provides up to a 10.1% average accuracy improvement and up to a 5.0% increase in the hard-level subset across five MLLMs, surpassing all top-notch related methods. Our code is available at https://github.com/AnonyNLP/NatLan.
@@ 1 Introduction
Figure 1: To address the failure of knowledge elicitation when directly answering in non-native language (Left), NatLan simulates the Positive Native Language Transfer of a human multilingual by utilizing two different MLLMs (Right). English meanings appear in <>absent<\ >< >. The roles involved in the process are displayed at the top.
Multilingual large language models (MLLMs) Brown et al. (2020); Achiam et al. (2023) have propelled the advancement of nearly all natural language processing tasks across various languages Xu et al. (2024); Li et al. (2024). However, it’s observed that MLLMs perform poorly on questions articulated in non-dominant languages, as depicted in Figure 1 (Left), failing to answer some questions that they could address when presented in their dominant language (i.e., the language with the highest proportion during training, such as English for Llama Touvron et al. (2023a), which accounts for over 70% of the tokens in the pretraining corpus). This indicates that MLLMs are unable to effectively apply the rich knowledge acquired from dominant language contexts when operating in non-dominant language contexts. Although this inability can be attributed to the insufficient training arising from the differing volumes Xue et al. (2021); ImaniGooghari et al. (2023) and quality Sitaram et al. (2023) of training data across various languages, constructing large-scale, high-quality data across all languages is extremely labor-intensive and not feasible.
@@ 5 Experiments
In the proposed NatLan, the Transferor must be capable of translating the content from the target language (Chinese) into the native language (English) as accurately and coherently as possible. Therefore, we selected the Qwen series MLLMs Bai et al. (2023) as Transferors, for their leading capabilities in Chinese comprehension among all MLLMs. We chose Qwen models with 4B, 7B, and 14B parameters to analyze the effects of Transferors with varying capabilities on NatLan in §5.5 and §5.6. Additionally, we selected a five representative MLLMs with strong English comprehension skills and the capability to understand Chinese to serve as Speakers. These include models from the Phi Abdin et al. (2024), Gemma Team et al. (2024), Mistral Jiang et al. (2023), and Llama Touvron et al. (2023b) series. For ease of joint deployment with the Transferor LLMs, all these Speaker LLMs possess a moderate parameter scale, ranging from 3.8B to 7B.
@@ A.2 Comparative Analysis of Chinese-to-English Translation Cases
The impact of the capital relocation on Palangkaraya is________\_\_\_\__ _ _ _.
A. It is conducive to alleviating housing shortages
B. It is conducive to alleviating traffic congestion
C. It is conducive to the control of environmental pollution
D. The improvement of infrastructure
@@ D. Ethanol
{CJK*}UTF8gbsn A. 在水中溶解性: $Na_{2}CO_{3}<NaHCO_{3}$
{CJK*}UTF8gbsn B. 与相同浓度酸反应的剧烈程度: $Na_{2}CO_{3}<NaHCO_{3}$
{CJK*}UTF8gbsn C. 热稳定性: $Na_{2}CO_{3}<NaHCO_{3}$
{CJK*}UTF8gbsn D. 二者间在一定条件下可相互转化